# TabICL: A Tabular Foundation Model for In-Context Learning on Large Data

### v1

@@ Appendix A Further Experiments
To obtain rough predictions of the time for a forward pass (training and inference) in TabICL and TabPFN2, we leverage the runtime complexity of the column- and row-wise attention modules. On a table with $n$ rows and $m$ columns, intra-column attention has a time complexity of $O(n^{2}m)$ since it is performed separately for each column, while intra-row attention has a time complexity of $O(nm^{2})$. Together, we obtain a complexity of $O(nm(n+m))$, where $nm$ is the number of cells of the table. We plot this quantity on the $x$-axis of Figure A.1 and fit models of the form
@@
with MSLE loss, that is, a MSE loss between the log-transformed time and the log-transformed prediction. To facilitate a simpler comparison, we fix $\gamma\coloneqq 0.8$ for both models since it yields a good fit. We suspect that $\gamma=1$ should be used asymptotically, but then more additional terms would be needed to obtain a good fit. In this model, the speedup of TabICL over TabPFNv2 for large datasets approaches $5$, while it is $1.4$ for small datasets.
Figure A.1: Training+inference times of TabICL and TabPFNv2. Each point represents the time on one dataset on an A100 GPU. We only use datasets with less than 30,000 samples because TabPFNv2 is applied with subsampling on larger datasets to avoid RAM overflow.
@@
Figure A.2 shows the scaling with the number of classes. The performance of TabICL deteriorates on datasets with three classes but not on datasets with more classes. It is unclear if this is really due to the number of classes or caused by a different characteristic present in the three-class datasets on the benchmark.
Figure A.3 shows the scaling with the number of features. These plots show that TabICL behaves well for large numbers of features even though it tokenizes entire rows in the middle of the network before seeing any labels.
Figure A.4 shows the dependence on the ratio of categorical to numerical variables. In general, the performance of TabICL and TabPFNv2 deteriorates somewhat in the presence of many categorical variables. However, it is notable that TabICL performs slightly better than TabPFNv2 on such datasets even though TabPFNv2 has a more sophisticated categorical feature generation in its prior.
Figure A.2: Dependency of the benchmark rank on the number of classes. Each point is the rank of one method on one dataset. Lower rank is better. The lines show the bootstrap median and 10% / 90% bootstrap confidence intervals of a piecewise linear fit.
Figure A.3: Dependency of the benchmark rank on the number of features. Each point is the rank of one method on one dataset. Lower rank is better. The lines show the bootstrap median and 10% / 90% bootstrap confidence intervals of a piecewise linear fit.
Figure A.4: Dependency of the benchmark rank on the ratio of categorical to overall features. Each point is the rank of one method on one dataset. Lower rank is better. The lines show the bootstrap median and 10% / 90% bootstrap confidence intervals of a piecewise linear fit.
Appendix B Synthetic Datasets for Pretraining
B.1 SCM prior with more activation functions
@@
    Here, all random parameters are drawn once per layer, and $\odot$ is an element-wise product. This is motivated by the fact that for fixed feature map $\phi$, the random function $f(x)=\phi(x)^{\top}\mathbf{z}$ is a Gaussian process with covariance kernel $k(x,x^{\prime})=\phi(x)^{\top}\phi(x^{\prime})$. The design of $\phi$ is inspired by random Fourier features (Rahimi & Recht, 2007). The randomly drawn exponent $-\exp(u)$ leads to different decays of coefficients with increasing frequency, which produces different levels of smoothness for the sampled function as shown in Figure B.1 (right) (Da Costa et al., 2023). This activation function applies standardization directly before the random function, such that the random rescaling has no effect.
@@
Figure B.1 visualizes the used activation functions. Figure B.2 shows datasets from the resulting prior.
Figure B.1: Activation functions from the SCM prior. Left: Non-random activation functions from the SCM prior without standardization and random rescaling. Activation functions are randomly flipped along the $x$- and/or $y$-axis to fully utilize the space in the plot. Right: Random instantiations of the random activation function from the SCM prior, including the automatic standardization of the inputs.
Figure B.2: Randomly generated 2D datasets from the SCM prior. The color corresponds to the class label.
B.2 Tree-based SCM prior
@@
Figure B.3 shows datasets from the tree SCM prior.
Figure B.3: Randomly generated 2D datasets from the tree-based SCM prior. The color corresponds to the class label.
Appendix C Rotary Positional Embedding
@@
$\theta_{i}$ $\coloneqq\frac{p}{10000^{2i/d}}$
@@
Appendix D Setup of TabICL
D.1 Pretraining details
@@
    $N_{\mathcal{B}}=4$ with a fixed size of 1,024 the first 100K steps.
@@
Each step comprises 512 datasets. In the first stage, all datasets contain an equal number of samples. In the second and third stages, datasets in each micro-batch have the same number of samples, but the number of samples varies between different micro-batches. We use Adam (Kingma & Ba, 2014) and clip the gradient norm to 1. The learning rate schedules for pretraining are shown in Figure 1(c).
(a) Cosine decay with warmup for stage 1
@@
Figure D.1: Learning rate schedules for 3 pretraining stages.
D.2 Memory-efficient inference
@@
It is important to note that the specific meanings of batch size and sequence length vary across different transformers, as outlined below:
@@
Table D.1: Notions of batch size and sequence length for different transformers
@@
In addition to adjusting the batch size, we also offload intermediate activations to the CPU or disk as needed to further alleviate GPU memory constraints. Figure D.2 illustrates the CPU and GPU memory consumption for a large dataset containing 100K samples and 500 features (80% training set and 20% test set). As shown, only 5GB of GPU memory and 25GB of CPU memory are utilized, making this a highly affordable computational setup.
Figure D.3 shows the CPU and GPU memory consumption for a larger dataset containing 500K samples and 500 features (80% training set and 20% test set). As depicted, the computation requires less than 14GB of GPU memory. The CPU memory usage reaches approximately 120GB, which, however, can be significantly reduced through optional disk offloading via memory mapping.
@@
Figure D.2: CPU and GPU memory consumption during inference for a dataset with 100K samples and 500 features. During dataset-wise ICL, there is only a single batch (i.e., a single dataset), which corresponds to a sharp spike in GPU usage at the final stage of computation.
@@
Figure D.3: CPU and GPU memory consumption during inference for a dataset with 500K samples and 500 features. The relatively high CPU memory consumption can be largely mitigated by optional disk offloading via memory mapping.
Appendix E Average Performance and Rankings
@@
It is important to note that accuracy is used as the objective metric for hyperparameter tuning in tabular methods.
@@
Figure E.1: Critical difference diagram for all 188 classification datasets ($\leq$ 10 classes).
@@
Figure E.2: Critical difference diagram for 120 binary classification datasets.
@@
Figure E.3: Critical difference diagram for 68 multi-class classification datasets ($\leq$ 10 classes).
@@
Figure E.4: Critical difference diagram for 132 small classification datasets ($\leq$ 10K samples).
@@
Figure E.5: Critical difference diagram for 56 large classification datasets ($>$ 10K samples).

### v2

@@ Appendix B Further Experiments
To obtain rough predictions of the time for a forward pass (training and inference) in TabICL and TabPFNv2, we leverage the runtime complexity of the column- and row-wise attention modules. As discussed in Section 3.5, for a table with $n$ rows and $m$ columns, the runtime complexity is $O(n^{2}m+nm^{2})=O(nm(n+m))$ for TabPFNv2, while it is $O(n^{2}+nm^{2})$ for TabICL. For simplicity, we use the term $nm(n+m)$ from the complexity of TabPFNv2, plot this quantity on the $x$-axis of Figure B.1, and fit models of the form
@@
with MSLE loss, that is, a MSE loss between the log-transformed time and the log-transformed prediction. To facilitate a simpler comparison, we fix $\gamma\coloneqq 0.8$ for both models since it yields a good fit. We suspect that $\gamma=1$ should be used asymptotically for TabPFNv2, but then more additional terms would be needed to obtain a good fit. In this model, the speedup of TabICL over TabPFNv2 for large datasets approaches $5$, while it is $1.4$ for small datasets.
Figure B.1: Training+inference times of TabICL and TabPFNv2. Each point represents the time on one dataset on an A100 GPU. We only use datasets with less than 30,000 samples because TabPFNv2 is applied with subsampling on larger datasets to avoid RAM overflow.
@@
Figure B.2 shows the scaling with the number of classes. The performance of TabICL deteriorates on datasets with three classes but not on datasets with more classes. It is unclear if this is really due to the number of classes or caused by a different characteristic present in the three-class datasets on the benchmark.
Figure B.3 shows the scaling with the number of features. These plots show that TabICL behaves well for large numbers of features even though it tokenizes entire rows in the middle of the network before seeing any labels.
Figure B.4 shows the dependence on the ratio of categorical to numerical variables. In general, the performance of TabICL and TabPFNv2 deteriorates somewhat in the presence of many categorical variables. However, it is notable that TabICL performs slightly better than TabPFNv2 on such datasets even though TabPFNv2 has a more sophisticated categorical feature generation in its prior.
Figure B.2: Dependency of the benchmark rank on the number of classes. Each point is the rank of one method on one dataset. Lower rank is better. The lines show the bootstrap median and 10% / 90% bootstrap confidence intervals of a piecewise linear fit.
Figure B.3: Dependency of the benchmark rank on the number of features. Each point is the rank of one method on one dataset. Lower rank is better. The lines show the bootstrap median and 10% / 90% bootstrap confidence intervals of a piecewise linear fit.
Figure B.4: Dependency of the benchmark rank on the ratio of categorical to overall features. Each point is the rank of one method on one dataset. Lower rank is better. The lines show the bootstrap median and 10% / 90% bootstrap confidence intervals of a piecewise linear fit.
Appendix C Synthetic Datasets for Pretraining
C.1 SCM prior with more activation functions
@@
    Here, all random parameters are drawn once per layer, and $\odot$ is an element-wise product. This is motivated by the fact that for fixed feature map $\phi$, the random function $f(x)=\phi(x)^{\top}\mathbf{z}$ is a Gaussian process with covariance kernel $k(x,x^{\prime})=\phi(x)^{\top}\phi(x^{\prime})$. The design of $\phi$ is inspired by random Fourier features (Rahimi & Recht, 2007). The randomly drawn exponent $-\exp(u)$ leads to different decays of coefficients with increasing frequency, which produces different levels of smoothness for the sampled function as shown in Figure C.1 (right) (Da Costa et al., 2023). This activation function applies standardization directly before the random function, such that the random rescaling has no effect.
@@
Figure C.1 visualizes the used activation functions. Figure C.2 shows datasets from the resulting prior.
Figure C.1: Activation functions from the SCM prior. Left: Non-random activation functions from the SCM prior without standardization and random rescaling. Activation functions are randomly flipped along the $x$- and/or $y$-axis to fully utilize the space in the plot. Right: Random instantiations of the random activation function from the SCM prior, including the automatic standardization of the inputs.
Figure C.2: Randomly generated 2D datasets from the SCM prior. The color corresponds to the class label.
C.2 Tree-based SCM prior
@@
Figure C.3 shows datasets from the tree SCM prior.
Figure C.3: Randomly generated 2D datasets from the tree-based SCM prior. The color corresponds to the class label.
Appendix D Rotary Positional Embedding
@@
$\theta_{i}$ $\coloneqq\frac{p}{100000^{2i/d}}$
@@
Appendix E Setup of TabICL
E.1 Pretraining details
@@
    $N_{\mathcal{B}}=4$ with a fixed size of 1,024 the first 160K steps.
@@
Each step comprises 512 datasets. In the first stage, all datasets contain an equal number of samples. In the second and third stages, datasets in each micro-batch have the same number of samples, but the number of samples varies between different micro-batches. We use Adam (Kingma & Ba, 2014) and clip the gradient norm to 1. The learning rate schedules for pretraining are shown in Figure E.1, including:
  • Cosine decay with restarts for stage 1
  • Polynomial decay for stage 2 and the learning rate is given by $(lr_{\text{init}}-lr_{\text{end}})\left(1-\text{step}/T\right)^{2}+lr_{\text{end}}$, where $lr_{\text{init}}=2e\text{-}5$, $lr_{\text{end}}=5e\text{-}6$, and the number of steps $T$ is 2,000.
  • Constant learning rate for stage 3
In addition, Figure E.2 shows critical difference diagrams across the three stages of curriculum learning. We observe that average rank of TabICL improves consistently from 11.4 (ninth place), to 7.46 (second), and finally to 6.95 (first). This provides strong empirical evidence for the effectiveness of curriculum learning.
(a) Cosine decay with restarts for stage 1
@@
Figure E.1: Learning rate schedules for 3 pretraining stages.
(a) Critical difference diagram (based on accuracy) after the stage 1 of curriculum learning
(b) Critical difference diagram (based on accuracy) after the stage 2 of curriculum learning
(c) Critical difference diagram (based on accuracy) after the stage 3 of curriculum learning
Figure E.2: Critical difference diagrams for different stages of curriculum learning
E.2 Memory-efficient inference
@@
It is important to note that the specific meanings of batch size and sequence length vary across different transformers, as outlined in Table E.1.
@@
Table E.1: Notions of batch size and sequence length for different transformers
@@
In addition to adjusting the batch size, we also offload intermediate activations to the CPU or disk as needed to further alleviate GPU memory constraints. Figure E.3 illustrates the CPU and GPU memory consumption for a large dataset containing 100K samples and 500 features (80% training set and 20% test set). As shown, only 5GB of GPU memory and 25GB of CPU memory are utilized, making this a highly affordable computational setup.
Figure E.4 shows the CPU and GPU memory consumption for a larger dataset containing 500K samples and 500 features (80% training set and 20% test set). As depicted, the computation requires less than 14GB of GPU memory. The CPU memory usage reaches approximately 120GB, which, however, can be significantly reduced through optional disk offloading via memory mapping.
@@
Figure E.3: CPU and GPU memory consumption during inference for a dataset with 100K samples and 500 features. During dataset-wise ICL, there is only a single batch (i.e., a single dataset), which corresponds to a sharp spike in GPU usage at the final stage of computation.
@@
Figure E.4: CPU and GPU memory consumption during inference for a dataset with 500K samples and 500 features. The relatively high CPU memory consumption can be largely mitigated by optional disk offloading via memory mapping.
E.3 Hierarchical Class-extension Strategy
Due to pretraining limitations, TabICL cannot natively handle classification problems with more than 10 classes. To address this, we employ hierarchical classification to extend the class scalability of TabICL. This involves building a multi-level classification tree where:
  1. 1.
    Each internal node splits the classes assigned to it into at most 10 child groups.
  2. 2.
    Recursion stops when a node contains $\leq 10$ classes, turning it into a leaf.
Since each split reduces the class count by at least a factor of 10, the tree depth $r=\Bigl{\lceil}\log_{10}k\Bigr{\rceil}$, where $k$ denotes the total number of classes.
Within the tree, each internal node predicts the probabilities of its child groups, while each leaf predicts the probabilities of the classes it contains. For any given class $c$, the unique path from the root to its containing leaf can be represented by a sequence of nodes:
$g_{0}(c)=\text{root},\quad g_{1}(c),\,g_{2}(c),\dots,\,g_{r-1}(c),\quad g_{r}(c)=c$
where $g_{i}(c)$ denotes the node at depth $i$ on the path leading to class $c$. Applying the probability chain rule along the path yields the overall posterior probability:
$p\!\bigl{(}y=c\mid x\bigr{)}\;=\;\prod_{i=1}^{r}p\!\bigl{(}g_{i}(c)\,\bigl{|}\,g_{i-1}(c),\,x,\,\mathcal{D}_{g_{i-1}(c)}\bigr{)}$ (7)
where $x$ is the given test sample, and $\mathcal{D}_{g_{i-1}(c)}$ represents the training data associated with the node $g_{i-1}(c)$. Note that the targets in $\mathcal{D}_{g_{i-1}(c)}$ are the child group indices for the node $g_{i-1}(c)$.
Appendix F Excluded Development Datasets of TabPFNv2
TabPFNv2 utilized several development datasets for hyperparameter tuning (e.g., learning rate, prior-related parameters), early stopping to prevent overfitting by monitoring performance on these datasets, and model selection to identify the best-performing model among multiple candidates. These development datasets are provided in the supplementary material of TabPFNv2 (Hollmann et al., 2025).
For a fair comparison, we excluded these development datasets from the main paper. In the table below, we report the performance of several high-performing methods on the development datasets included in the TALENT benchmark:
Table F.1: Performance on 15 development datasets of TabPFNv2
Dataset TabICL TabPFNv2 TabR ModernNCA RealMLP TabM CatBoost
Acc. AUC LogL Acc. AUC LogL Acc. AUC LogL Acc. AUC LogL Acc. AUC LogL Acc. AUC LogL Acc. AUC LogL
allbp 0.973 0.975 0.081 0.978 0.978 0.070 0.970 0.915 0.150 0.971 0.969 0.093 0.971 0.845 0.145 0.974 0.910 0.252 0.978 0.942 0.070
baseball 0.933 0.963 0.160 0.940 0.969 0.155 0.947 0.965 0.171 0.927 0.951 0.207 0.937 0.899 0.248 0.943 0.964 0.155 0.939 0.965 0.157
delta_ailerons 0.952 0.981 0.161 0.954 0.982 0.159 0.948 0.978 0.175 0.949 0.980 0.165 0.945 0.976 0.184 0.950 0.979 0.172 0.948 0.979 0.164
eye_movements 0.639 0.822 0.757 0.788 0.930 0.488 0.970 0.996 0.184 0.986 0.998 0.090 0.767 0.900 0.626 0.994 1.000 0.043 0.719 0.883 0.623
eye_movements_bin 0.588 0.648 0.655 0.678 0.751 0.597 0.659 0.732 1.344 0.909 0.974 0.207 0.596 0.639 2.244 0.593 0.628 0.668 0.615 0.672 0.642
heloc 0.720 0.799 0.545 0.726 0.803 0.543 0.722 0.795 0.555 0.721 0.797 0.547 0.722 0.798 0.551 0.718 0.789 0.557 0.723 0.799 0.544
hill-valley 0.785 0.917 0.518 0.983 0.992 0.098 0.700 0.732 0.571 0.781 0.804 1.455 0.796 0.852 0.547 0.512 0.529 0.703 0.517 0.515 0.693
JapaneseVowels 0.997 1.000 0.011 0.997 1.000 0.008 0.996 1.000 0.016 0.996 1.000 0.013 0.994 1.000 0.113 0.993 1.000 0.051 0.983 1.000 0.048
led24 0.731 0.953 0.812 0.732 0.956 0.780 0.738 0.949 0.885 0.732 0.955 0.973 0.734 0.954 0.827 0.734 0.958 0.793 0.735 0.956 0.962
national-longitudinal 0.998 1.000 0.005 1.000 1.000 0.000 0.994 1.000 0.029 0.999 1.000 0.003 0.995 0.996 0.064 1.000 1.000 0.000 1.000 1.000 0.012
page-blocks 0.976 0.993 0.092 0.975 0.993 0.092 0.968 0.987 0.127 0.967 0.989 0.129 0.971 0.980 0.162 0.968 0.985 0.127 0.967 0.991 0.110
ringnorm 0.980 0.997 0.066 0.980 0.997 0.055 0.980 0.995 0.075 0.980 0.996 0.072 0.976 0.996 0.079 0.980 0.997 0.062 0.971 0.995 0.083
rl 0.763 0.853 0.472 0.860 0.937 0.322 0.878 0.934 0.334 0.837 0.916 0.378 0.747 0.824 0.548 0.787 0.869 0.552 0.786 0.863 0.463
thyroid-ann 0.987 0.999 0.028 0.995 1.000 0.017 0.986 0.996 0.068 0.987 0.996 0.088 0.992 0.998 0.087 0.992 1.000 0.024 0.995 1.000 0.012
waveform-5000 0.872 0.978 0.259 0.865 0.977 0.263 0.860 0.975 0.310 0.862 0.975 0.278 0.867 0.975 0.291 0.862 0.976 0.279 0.860 0.973 0.297
Appendix G Random Forest Extension of TabPFNv2 for Large Datasets
Hollmann et al. (2025) proposed an extension of TabPFNv2 using random forest (RF) to enhance its performance on large datasets. During training, the extension recursively partitions the feature space using standard decision tree splitting criteria. At inference time, instead of relying on simple majority voting at the leaves, it fits TabPFNv2 to the subset of training data that reaches each leaf. This extension not only leverages the inductive biases of tree-based models but also reduces the number of samples processed by TabPFNv2, thereby mitigating its scalability limitations.
The ensemble prediction of this RF extension involves two main steps: (1) routing each test sample through all trees to their corresponding leaf nodes, and (2) aggregating the predictions from TabPFNv2 at these leaves by either averaging probabilities or logits.
A key innovation of this RF extension is the adaptive pruning mechanism that dynamically determines whether to use parent node predictions or leaf node predictions based on validation performance. This prevents overfitting and optimizes the tree depth for each region of the feature space.
This extension is not exclusive to TabPFNv2 and can also be applied to other tabular foundation models like TabICL. We evaluated this RF extension with adaptive pruning enabled for both TabPFNv2 and TabICL on large datasets with more than 10K samples. Additionally, we set the ensemble size for both TabPFNv2 and TabICL to 4, considering that random forest already ensembles multiple decision trees. As shown in Figure G.1, the RF extension significantly improves the performance of both TabPFNv2 and TabICL.
Figure G.1: Critical difference diagram (based on accuracy) on 53 large datasets with more than 10K samples. For TabICL-RF and TabPFNv2-RF, the ”RF” suffix indicates the use of the random forest extension.
Appendix H Effect of Ensemble Size on TabPFNv2 and TabICL
Both TabPFNv2 and TabICL break column-order invariance to mitigate the learning collapse problem, as described in Section 3.3. To approximately restore permutation invariance, they aggregate predictions across multiple column permutations, while also using different preprocessors and shuffling class labels. We analyze the effect of ensemble size on their performance, as shown in Figure H.1. Our findings indicate that TabPFNv2 shows a clear improvement even with a small ensemble size ($\leq 4$), whereas TabICL requires a larger ensemble size ($\geq 8$) for a noticeable performance gain.
Overall, TabPFNv2 benefits more from ensembling than TabICL. This could be attributed to :
  • TabPFNv2 employs more sophisticated preprocessing operations compared to TabICL.
  • TabPFNv2 benefits more from column permutations inherently. Unlike TabICL, TabPFNv2 never collapses the column dimension. This means permuting columns could have a more pronounced impact on the outputs of TabPFNv2, which reduces the correlation between ensemble members of TabPFNv2, thereby improving the effectiveness of ensembling.
(a) TabPFNv2
(b) TabICL
Figure H.1: Effect of ensemble size on TabPFNv2 and TabICL. The figures show the relative accuracy improvements of different ensemble sizes over no ensembling across 171 datasets with at most 10 classes.
Appendix I Average Performance and Rankings
@@
It is important to note that accuracy is used as the objective metric for hyperparameter tuning in tabular methods. Moreover, LoCalPFN (Thomas et al., 2024) and TuneTables (Feuer et al., 2024) are only included in the accuracy-based rankings.
@@
Figure I.1: Critical difference diagram for all 171 classification datasets ($\leq$ 10 classes).
@@
Figure I.2: Critical difference diagram for 112 binary classification datasets.
@@
Figure I.3: Critical difference diagram for 59 multi-class classification datasets ($\leq$ 10 classes).
@@
Figure I.4: Critical difference diagram for 116 small classification datasets ($\leq$ 10K samples).
@@
Figure I.5: Critical difference diagram for 53 large classification datasets ($>$ 10K samples).